OpenStudy (anonymous):

The polynomial 1-x+x ^{2}-x ^{3}+...-x ^{15}+x ^{16}-x ^{17} can be written as a polynomial in y=x+1. Find the coefficient of y ^{2}

OpenStudy (amistre64):

i cant make head nor hair of it

myininaya (myininaya):

So I guess we should replace x with y-1

myininaya (myininaya):

\[1-(y-1)+(y-1)^2-(y-1)^3+(y-1)^4 \cdots \] we really don't need to right this any further all of them will be larger degree than y^2

myininaya (myininaya):

write*

myininaya (myininaya):

that is how i see it anyways

OpenStudy (zarkon):

you could look at the Taylor polynomial

OpenStudy (amistre64):

doesnt (y-1)^3 expand with a y^2 in it?

OpenStudy (amistre64):

y^3, 3y^2, 3y, 1

OpenStudy (amistre64):

expanding all those terms would create y^2 s

myininaya (myininaya):

oh yeah you are right

myininaya (myininaya):

that wouldn't be efficient lol

myininaya (myininaya):

so we should look at zarkon's way

OpenStudy (amistre64):

well, if its the y^2 coeffs; its the pascal triangle diag then

OpenStudy (amistre64):

2,1,0 ^ 1 3 6 ...

OpenStudy (amistre64):

assuming myins interp is good :)

OpenStudy (amistre64):

..... the L of the pascal .... thats equal to the next ones entry ....

OpenStudy (amistre64):



OpenStudy (zarkon):

you want a polly that looks like this \[\sum_{k=0}^{17}a_ky^k\] there \(y=x+1\)

OpenStudy (zarkon):

*where y=x+1

OpenStudy (amistre64):

1 1* 1 1 2* 1 1 3 3* 1 1 4 6 4* 1 1510 (10) 5 1

OpenStudy (amistre64):

but i think my idea has + and - signs to deal with still

OpenStudy (zarkon):

I get 816 as the answer

OpenStudy (amistre64):

\[{{18}\choose {3}}=816.too\]

OpenStudy (asnaseer):

I notice that the person you guys are trying to assist has not said a word yet. @kwenisha - does all this make any sense to you? have you attempted it yourself using a different method? any thoughts?

OpenStudy (zarkon):

using the Taylor polly amounts to computing \[\frac{\displaystyle\sum_{k=1}^{17}k(k-1)}{2}\]

OpenStudy (anonymous):

@Asnaseer, I am actually trying understand it while working with my partner... I apologize if I come off as just trying to get the answer and not trying to understand it...

OpenStudy (asnaseer):

thats fine @kwenisha - I was just worried in case the experts above left you dazed. :)